Solve the following equation for x.
H=1/4(x-2y)

Answer by Alan3354(69443)

Solve the following equation for x.
H=1/4(x-2y)
If you mean H=(1/4)(x-2y)
Multiply by 4
4H = x-2y
Add 2y
4H + 2y = x
x = 4H + 2y